OpenStudy (anonymous):

(3/2)^4 = evaluate to improper fraction

OpenStudy (goldphenoix):

\[\large \large (3/2)^4 = \frac{ 3 }{ 2 }\times \frac{ 3 }{ 2 }\times\frac{ 3 }{ 2 }\times\frac{ 3 }{ 2 }\] To multiply fractions, you just multiply the numerator and the denominator. 3 * 3 *3 *3 = 81 2 * 2 *2 *2 = 16. So it would look like: \[\large \large \frac{ 81 }{ 16 }\]